After a sailboat had been on the water for 3h, a change in the wind direction reduced the average speed of the boat by 5 mph. The entire distance sailed was 57 mi. The total time spent sailing was 6 h. How far did the sailboat travel in the first 3 h?
Answer by stanbon(75887)

After a sailboat had been on the water for 3h, a change in the wind direction reduced the average speed of the boat by 5 mph. The entire distance sailed was 57 mi. The total time spent sailing was 6 h. How far did the sailboat travel in the first 3 h?
-----------------------
1st part DATA:
time = 3 hr ; rate = x mph ; distance = rt = 3x miles
-------------
2nd part DATA:
time = 3 hr ;rate = x-5 mph ; distance = rt = 5x-25 miles
-------------
Equation:
distance + distance = 57
3x + 5x-25 = 57
8x = 82
x = 10 1/4 mph
----------------------
Distance traveled in 1st 3 hr. = (10 1/4)*3 = 30 3/4 miles
==================================================================
Cheers,
Stan H.
Reply to stanbon@comcast.net